# Laura Barros        Aiko Schmeißer

### v1

@@ Abstract
The supply of politicians affects the quality of democratic institutions. Yet, little is known about how economic trade-offs motivate individuals to enter political life. This paper investigates how experiencing a job loss affects individuals’ decision to join politics and studies the consequences of layoffs for the selection of politicians. Using administrative data on the universe of formal employees, party members, and local political candidates in Brazil, and relying on mass layoffs for causal identification, we show that job loss increases the probability of joining a political party and running for local office. Moreover, we document that layoff-induced candidates are positively selected in various competence measures, suggesting that economic shocks may improve the quality of politicians. Heterogeneity analyses reveal a more pronounced increase in candidacies for laid-off individuals with higher financial incentives from office holding and larger predicted income losses. In addition, we find that being eligible for unemployment benefits after job loss also increases party memberships and candidacies. These results are consistent with the reduction in private-sector opportunity costs and the increased time resources explaining the rise in political entry.
@@ 1 Introduction
In this paper, we study how experiencing a job loss affects individuals’ decision to enter politics and examine the consequences of these economic shocks for the quality of political candidates. Using rich administrative data from Brazil, we focus on party memberships and candidacies for local councils, the lowest-level positions in Brazilian politics. We rely on individual-level information on the universe of party members and local candidates between 2000 and 2020, which we match to employer-employee data on all formally employed workers between 2004 and 2018. The comprehensiveness of the data allows us to investigate effect heterogeneity along individual, party, and institutional characteristics and explore the consequences of eligibility for unemployment benefits after the job loss. In addition, we use various competence measures to document the implications of job loss for the selection of those who enter politics. Taken together, our results document novel insights on the mechanisms behind individual decisions to engage with politics following a layoff, as well as the consequences of these economic shocks for the supply of politicians.
@@
To estimate the effects of job loss on the probability of entering politics, we exploit firm-level mass layoffs as plausibly exogenous variation for individual-level job loss. We compare individuals laid off in a mass layoff firm to a matched control group of similar workers from other firms who were not dismissed in the same year. The richness of our data allows us to exactly match treated and control workers on several characteristics, such as gender, education, age, wage, tenure, and firm size, and to control for fine-grained local- and industry-specific shocks that may jointly affect mass layoffs and political behavior. Importantly, accounting for changes in the local political environment ensures that our results capture individual supply responses and are not influenced by changes in the demand for politicians. We use a difference-in-differences (DiD) design to estimate dynamic treatment effects for up to three election cycles (12 years) after job loss. The underlying parallel-trends assumption is supported by the absence of differential trends in political outcomes between treated and control individuals before the layoff, and we also provide evidence that selection into (mass-)layoffs does not drive our results. [^1] Our empirical strategy closely follows recent papers studying the effects of mass layoffs in Brazil on crime (Britto et al., 2022b), domestic violence (Bhalotra et al., 2021), health (Amorim et al., 2023), and children’s education (Britto et al., 2022a).
@@
Our main results show that experiencing a job loss increases the probability of joining a party by more than 12% compared to the control mean. This increase in party membership is long-lasting, starting in the election cycle after job loss and remaining statistically significant in the two subsequent election cycles. For the probability of running for local office, the effect increases over time, ranging from 6% in the first election cycle after layoff up to 28% in the two subsequent election cycles. The dynamic effect observed for candidacies is consistent with the fact that local councilors have no term limit and often re-run for office after their first candidacy. Additionally, the process of deciding to become a candidate and preparing an election campaign requires a substantial amount of time and planning, consistent with the delayed effects we estimate.
@@
The positive effects of layoffs on political engagement are remarkably widespread. First, in robustness tests, we show that the effects do not differ strongly by the size of a mass layoff, mitigating concerns regarding external validity to regular layoffs. Second, we examine heterogeneity across a range of individual covariates (education, age, gender, wage, tenure) and local characteristics (average income, unemployment rates, inequality, homicide rates, left vs. center-right presidential vote shares). We find significant positive effects on party memberships and candidacies in all the considered individual and local subsamples. Third, we show that the effects of job loss extend beyond the realms of party memberships and councilor candidacies, encompassing various facets of political participation. Specifically, we also find positive effects on the likelihood of running for mayor, donating to a political candidate, and working for a political campaign.
@@
What explains the positive effects of job loss on political entry? Economic theories of political selection model the decision to run for office as a trade-off between the expected benefits from holding office and the costs of running (Black, 1972; Caselli and Morelli, 2004; Messner and Polborn, 2004). A layoff decreases the opportunity costs of entering politics by lowering the private-sector outside options and increasing the time available for political engagement. The size and granularity of our data enable us to examine heterogeneity in treatment effects and to test for the role of the opportunity cost channel in driving the increase in political entry after job loss. We start by exploiting local variation in councilor salaries – that depend on the size of Brazilian municipalities (Ferraz and Finan, 2011) – and individual variation in workers’ predicted income losses. The former captures the financial attractiveness of a political career, while the latter captures the severity of the drop in private-sector opportunities predicted from individual observables and labor market characteristics. We find larger layoff effects on candidacies in municipalities with higher wage caps, with no corresponding effects for memberships. Consistently, the increase in candidacies is highest among individuals who are predicted to face the greatest income loss due to a layoff. Together, these results show the importance of monetary incentives for individuals’ decision to enter politics following a job loss.
@@
In order to explore the role of time resources, we analyze the effects of eligibility for benefits from unemployment insurance (UI). During benefit receipt, individuals have lower incentives to find new employment and thus more time available for engaging with politics. We conduct a regression discontinuity analysis exploiting the fact that benefit eligibility varies discontinuously with the time since the last layoff (a minimum of 16 months is required to be able to file a new UI claim). We find that benefit eligibility reduces employment and increases the likelihood of becoming a party member or running for local office. The effects are larger for individuals with higher education, amounting to 6.2% increase in memberships and a 13.0% increase in candidacies (relative to the control mean) in the two election cycles after the layoff. These results are robust to varying bandwidth choices, polynomial specifications, and permutation tests and point towards the importance of time resources for competent individuals to enter politics after layoffs.
@@
To test for these mobilization mechanisms, we disaggregate memberships and candidacies in different parties. We start by examining ideological differences, classifying parties as left, center, and right. We find that job loss increases memberships and candidacies in parties across the ideological spectrum. [^2] The only notable heterogeneity is a larger positive effect on right-wing party memberships in the 2020 cycle, the term of far-right president Jair Bolsonaro. We also disentangle the party of the local incumbent mayor and other non-incumbent parties. The results reveal positive participation effects for memberships and candidacies in both incumbent and non-incumbent parties. Finally, we look at the role of union affiliation by identifying parties historically aligned with labor unions. We document that following layoffs, there is an increase in individuals’ likelihood of becoming members and running for office both at union-affiliated and non-union-affiliated parties. Overall, the heterogeneity analyses indicate that incumbent, center, and non-union-affiliated parties also experience increased memberships and candidacies following layoffs, suggesting that mobilization and politicization cannot fully explain the rise in participation.
@@
A crucial question that follows is how layoffs affect the selection of politicians. Our results suggest that individuals’ decision to engage with politics following layoffs is at least partly explained by economic rather than purely ideological motives. If economic incentives vary across individuals’ skill levels, economic shocks may also impact the average competence of politicians and, thereby, the functioning of the political system. To study the implications for the selection of political participants, we provide different pieces of evidence. First, we compare the characteristics of treated and control individuals who become party members and candidates after the layoff period. Given that we match individuals along observable characteristics, treated and control groups have identical pre-layoff characteristics in the full sample. Thus, any observable differences in pre-layoff characteristics of treated and control workers who engage with politics after the layoff must stem from the heterogeneous effects of job loss across individuals. We document that treated members and candidates are positively selected in terms of education, previous earnings, and political experience as compared to those who do not experience a layoff. We find the same results when considering a measure of relative earnings capacity that is calculated as the residual from a Mincer earnings regression and more likely reflects personal ability. Second, to study aggregate implications, we estimate the effects of municipal employment rates on the average education of local councilors. Exploiting local labor demand shocks from a Bartik (1991) instrument, we find that lower employment rates increase the average years of education not only among those individuals who run for office but also among those who are ultimately elected into councilor office. Overall, these results suggest that economic downturns can lead to positive political selection.
@@
Our paper contributes to two main strands of research. First, we speak to the literature that studies the determinants of political selection. [^3] See Dal Bó and Finan (2018) and Gulzar (2021) for recent reviews of this literature. In the tradition of the citizen-candidate framework (Osborne and Slivinski, 1996; Besley and Coate, 1997), this literature highlights the importance of endogenous entry into politics. Individuals’ decision to run for office is modeled as a cost-benefit calculation, where the expected benefits of holding office are weighed against the costs of running. On the benefit side, Ferraz and Finan (2011) and Gagliarducci and Nannicini (2013) document that increasing politicians’ salaries attracts higher-quality candidates and improves the performance of incumbent politicians. In addition to salaries, holding a public office may lead to other non-salary earnings while in office (whether legal or not) and enhance private sector opportunities after leaving office (Eggers and Hainmueller, 2009; Fisman et al., 2014). On the cost side, running for office may involve not only monetary, time, and opportunity costs, but also psychological and family burdens (Hall, 2019). Beyond individual factors, political selection is shown to be affected by institutional constraints, such as electoral rules (Arora, 2022), gender quotas (Besley et al., 2017), and campaign spending limits (Avis et al., 2022). Our primary contribution to this line of research is to provide evidence on how economic trade-offs motivate individuals to enter politics. Specifically, we document how an adverse shock in private-sector opportunities increases individuals’ likelihood of starting a political career. The effects increase with several measures of individual competence, suggesting that layoffs can improve candidate quality. In addition, we show that unemployment benefits, by changing individuals’ time constraints after layoffs, impact political selection, resulting in an increased entry of better-educated individuals into politics. To the best of our knowledge, we are the first to study the impact of layoffs on political entry using large-scale individual-level data, which allows us to tackle the rare-event problem that researchers typically face in the study of political selection.
@@
Second, we add to the literature on the political consequences of economic shocks. A growing number of papers documents that voting results respond to macro shocks, such as import competition (Autor et al., 2020; Dippel et al., 2022), economic crisis (Barros and Santos Silva, 2019), automation (Anelli et al., 2019), and austerity reforms (Fetzer, 2019). Similarly, at the individual level, many studies document that changes in economic conditions affect individuals’ policy preferences, including support for redistribution and welfare provision (e.g. Margalit, 2013; Martén, 2019; Ahlquist et al., 2020; Ballard-Rosa et al., 2021). In terms of political participation, the literature has focused mostly on turnout and has found mixed results, ranging from turnout increasing with local unemployment rates (Charles and Stephens, 2013; Burden and Wichowsky, 2014), decreasing after individual job loss (Emmenegger et al., 2017; Östermann and Lindgren, 2023), or not being affected at all by income changes (Jungkunz and Marx, 2022; Geys and Sørensen, 2024). Our contribution to this strand of literature is to show that economic shocks not only affect the demand of voters, but also impact the supply side of politics. In particular, we document how layoffs increase individuals’ propensity to enter the political system, either as members of a political party or as candidates in local elections. In addition, our mass-layoff identification strategy, together with the size and granularity of our data, allows us to exploit exogenous changes in personal economic conditions and study the mechanisms behind individuals’ decisions to engage with politics after a layoff.
@@
This paper is structured as follows. Section 2 provides background information on party members and local council candidates in Brazil, and discusses theoretical predictions about the role of opportunity costs for political entry. In Section 3, we present the datasets used in the analyses, and in Section 4, we outline our empirical strategy. Section 5 presents the main results and heterogeneity analyses for the effect of job loss. In Section 6, we investigate the role of UI benefit eligibility. Finally, Section 7 examines potential implications for political selection and Section 8 concludes.
@@ 2 Institutional and Theoretical Background
In Brazil, a significant portion of the population is engaged in politics. As mandated by the 1988 Brazilian constitution, voting is compulsory for all literate citizens aged 18 to 69.444The average turnout rate in Federal Elections in the last 20 years was about 80% (International Institute for Democracy and Electoral Assistance, 2024). Apart from voting, more than 16 million Brazilians were registered as members of a political party in 2020. This corresponds to about 11% of the electorate, one of the highest shares across democracies. In addition, between 2000 and 2020, about 1.7 million individuals have run for office in local elections in Brazil. In this section, we describe the institutional environment in which individuals decide about their involvement as party members or local candidates and present a theoretical framework that illustrates the role of opportunity costs in this decision.
@@ 2.1 Party members
Membership recruitment in Brazil is predominantly a local phenomenon, fueled by internal competition at parties’ municipal conventions. Local elites, aiming to secure the party’s nomination, attract supporters by enlisting them in the party to build a loyal voting base at the convention and signal their potential as candidates (Sells, 2020; Frey, 2022). Supporters who become members of a winning political party are rewarded with employment opportunities within the government bureaucracy (Brollo et al., 2017; Barbosa and Ferreira, 2023). The importance of members for local politics is evident from the cyclical pattern of new affiliations: as we show in Appendix Figure A.1, parties significantly expand their ranks in the year before local elections.

### v3

@@ Abstract
The supply of politicians affects the quality of democratic institutions. Yet little is known about the economic motivations that drive individuals into politics. This paper examines how experiencing a job loss affects individuals’ decisions to enter political life and explores its implications for political selection. Using highly granular administrative data linking individual records of political participation with comprehensive employer-employee data for all formal workers in Brazil, and leveraging mass layoffs for causal identification, we find that job loss significantly increases the likelihood of joining a political party and running for local office. Layoff-induced candidates are positively selected on various competence measures, suggesting that economic shocks improve the quality of political entrants. Further, we observe that the increase in candidacies is more pronounced among laid-off individuals with greater financial incentives from holding office and higher predicted income losses. Finally, using a regression discontinuity design, we find that eligibility for unemployment benefits increases the likelihood of becoming a party member and running for local office. These results are consistent with the reduction in private-sector opportunity costs and the increased time resources explaining the rise in political entry.
@@ 1 Introduction
In this paper, we study how experiencing a job loss affects individuals’ decision to enter politics and examine the consequences of these economic shocks for the quality of political candidates. Using rich administrative data from Brazil, we focus on party memberships and candidacies for local councils, the lowest-level positions in Brazilian politics. We rely on individual-level information on the universe of party members and local candidates between 2000 and 2020, which we match to employer-employee data on all formally employed workers between 2004 and 2018. The comprehensiveness of the data allows us to investigate effect heterogeneity along individual, local, and party characteristics and explore the consequences of eligibility for unemployment benefits after the job loss. In addition, we use various competence measures to document the implications of job loss for the selection of those who enter politics. Taken together, our results document novel insights into the mechanisms behind individual decisions to engage with politics following displacement from the private labor market, as well as the consequences of such economic shocks for the supply of politicians.
@@
To estimate the effects of job loss on the probability of entering politics, we exploit firm-level mass layoffs as plausibly exogenous variation for individual-level job loss. We compare individuals laid off in a mass layoff firm to a matched control group of similar workers from other firms who were not dismissed in the same year. The richness of our data allows us to exactly match treated and control workers on several characteristics, such as gender, education, age, wage, tenure, and firm size, and to control for fine-grained local- and industry-specific shocks that may jointly affect mass layoffs and political behavior. Importantly, accounting for changes in the local political environment ensures that our results capture individual supply responses and are not influenced by changes in the demand for politicians. We use a difference-in-differences (DiD) design to estimate dynamic treatment effects for up to three election cycles (12 years) after job loss. The underlying parallel-trends assumption is supported by the absence of differential trends in political outcomes between treated and control individuals before the layoff, and we also provide evidence that selection into (mass-)layoffs does not drive our results. [^1] Our empirical strategy closely follows recent papers studying the effects of mass layoffs in Brazil on crime (Britto et al., 2022b), domestic violence (Bhalotra et al., 2025), health (Amorim et al., 2025), and children’s education (Britto et al., 2022a).
@@
Our main results show that experiencing a job loss increases the probability of joining a party by more than 12% compared to the control mean. This increase in party membership is long-lasting, starting in the election cycle after job loss and remaining statistically significant in the two subsequent election cycles. For the probability of running for local office, the effect increases over time, ranging from 6% in the first election cycle after layoff up to 28% in the two subsequent election cycles. The delayed increase in candidacies likely reflects the time needed to plan a campaign and the absence of term limits for local councilors. The persistence of effects is also consistent with the long-term scarring effects of job displacement, as we show that earnings in the formal labor market remain durably lower after the layoff.
@@
The positive effects of layoffs on political engagement are remarkably widespread. First, in robustness tests, we show that the effects do not differ strongly by the size of a mass layoff, mitigating concerns regarding external validity to regular layoffs. Second, we examine heterogeneity across a range of individual covariates (education, age, gender) and local characteristics (average income, unemployment rates, inequality, homicide rates, left vs. center-right presidential vote shares). We find significant positive effects on party memberships and candidacies in all the considered individual and local subsamples. Third, we show that the effects of job loss extend beyond the realms of party memberships and councilor candidacies, encompassing various facets of political participation. Specifically, we also find positive effects on the likelihood of running for mayor, donating to a political candidate, and working for a political campaign.
@@
What explains the positive effects of job loss on political entry? Economic theories of political selection model the decision to run for office as a trade-off between the expected benefits from holding office and the costs of running (Black, 1972; Caselli and Morelli, 2004; Messner and Polborn, 2004). A layoff decreases the opportunity costs of entering politics by lowering the private-sector outside options and increasing the time available for political engagement. The size and granularity of our data enable us to examine heterogeneity in treatment effects and to test for the role of the opportunity cost channel in driving the increase in political entry after job loss. We start by exploiting individual variation in workers’ predicted income losses from the layoff and local variation in councilor salaries which depend on the size of Brazilian municipalities (Ferraz and Finan, 2011). The former captures the severity of the drop in private-sector opportunities predicted from individual observables and labor market characteristics, while the latter captures the financial attractiveness of a political career. We find larger layoff effects on candidacies among individuals who are predicted to face the greatest income loss due to a layoff. Consistently, the increase in candidacies is highest in municipalities with higher wage caps, with no corresponding effects for memberships. Together, these results show the importance of monetary incentives for individuals’ decision to enter politics following a job loss.
@@
In order to explore the role of time resources, we analyze the effects of eligibility for benefits from unemployment insurance (UI). During benefit receipt, individuals have lower incentives to find new employment and thus more time available for engaging with politics. We conduct a regression discontinuity analysis exploiting the fact that benefit eligibility varies discontinuously with the time since the last layoff (a minimum of 16 months is required to be able to file a new UI claim). We find that benefit eligibility reduces employment and increases the likelihood of becoming a party member or running for local office. The effects are larger for individuals with higher education, amounting to 6.2% increase in memberships and a 13.0% increase in candidacies (relative to the control mean) in the two election cycles after the layoff. These results are robust to varying bandwidth choices, polynomial specifications, and permutation tests, and point towards the importance of time resources in enabling more educated individuals to enter politics after layoffs.
@@
To test for these mobilization mechanisms, we disaggregate memberships and candidacies in different parties. We start by examining ideological differences, classifying parties as left, center, and right. We find that job loss increases memberships and candidacies in parties across the ideological spectrum. [^2] The only notable heterogeneity is a larger positive effect on right-wing party memberships in the 2020 cycle, the term of far-right president Jair Bolsonaro. We also disentangle the party of the local incumbent mayor and other non-incumbent parties. The results reveal positive participation effects for memberships and candidacies in both incumbent and non-incumbent parties. Finally, we look at the role of union affiliation by identifying parties historically aligned with labor unions. We document that following layoffs, there is an increase in individuals’ likelihood of becoming members and running for office both at union-affiliated and non-union-affiliated parties. Overall, the heterogeneity analyses indicate that incumbent, center and non-union-affiliated parties also experience increased memberships and candidacies following layoffs, suggesting that mobilization and politicization cannot fully explain the rise in political participation.
@@
A crucial question that follows is how layoffs affect the selection of politicians. Our results suggest that individuals’ decisions to engage with politics following layoffs are at least partly driven by economic rather than purely ideological motives. If economic incentives vary across individuals’ skill levels, economic shocks may also influence the average competence of politicians and, in turn, the functioning of the political system. To study the implications for the selection of political participants, we provide two pieces of evidence. First, we compare the characteristics of individuals who become party members and candidates due to the layoff (compliers) with those who would engage with politics regardless (always-takers) and those who would not engage with politics under any circumstances (never-takers). We find that compliers are positively selected in terms of education, previous wages, and AKM person wage effects, compared to always-takers and never-takers. For candidates, we also document that compliers have higher levels of education relative to the universe of candidates. Importantly, this does not come at the expense of political representation, as our results indicate that individuals who decide to run for office due to the layoff are younger and more likely to be female in comparison to always-takers and the population of candidates. Second, to study aggregate implications, we estimate the effects of municipal employment rates on the average education, gender, and age of local councilors. Exploiting local labor demand shocks from a Bartik (1991) instrument, we find that lower employment rates increase the average years of education not only among those individuals who run for office but also among those who are ultimately elected as councilors. Additionally, we observe that lower employment rates lead to a higher share of female and young councilors. Overall, these results suggest that economic downturns can lead to more positive political selection without compromising political representation.
@@
Our paper contributes to two main strands of research. First, we speak to the literature that studies the determinants of political selection. [^3] See Dal Bó and Finan (2018) and Gulzar (2021) for recent reviews of this literature. In the tradition of the citizen-candidate framework (Osborne and Slivinski, 1996; Besley and Coate, 1997), this literature highlights the importance of endogenous entry into politics. Individuals’ decision to run for office is modeled as a cost-benefit calculation, where the expected benefits of holding office are weighed against the costs of running. On the benefit side, Ferraz and Finan (2011) and Gagliarducci and Nannicini (2013) document that increasing politicians’ salaries attracts higher-quality candidates and improves the performance of incumbent politicians. In addition to salaries, holding a public office may lead to other non-salary earnings while in office (whether legal or not) and enhance private sector opportunities after leaving office (Eggers and Hainmueller, 2009; Fisman et al., 2014). On the cost side, running for office may involve not only monetary, time, and opportunity costs, but also psychological and family burdens (Hall, 2019). Using detailed information on the social background of politicians in Sweden, Dal Bó et al. (2017) document that Swedish politicians are positively selected along various socio-demographic characteristics. In contrast, Britto et al. (2024) show that in Brazil, individuals with a criminal background are more likely to run for and win political office than the general population. Beyond individual determinants, political selection is also shaped by institutional factors, such as electoral and fiscal rules (Arora, 2022; Gamalerio and Trombetta, 2025), gender quotas (Besley et al., 2017), campaign spending limits (Avis et al., 2022), and political training programs (Dal Bó et al., 2024).
Our contribution to this line of research is to provide evidence on how economic trade-offs motivate individuals to enter politics. Specifically, we document how an adverse shock in private-sector opportunities increases individuals’ likelihood of starting a political career. The effects increase with several measures of individual competence, suggesting that layoffs can improve candidate quality. In addition, we show that unemployment benefits, by changing individuals’ time constraints after layoffs, impact political selection, resulting in an increased entry of better-educated individuals into politics. To the best of our knowledge, we are the first to study the impact of layoffs on political entry using large-scale individual-level data, which allows us to tackle the rare-event problem that researchers typically face in the study of political selection.
@@
Second, we add to the literature on the political consequences of economic shocks. A growing number of papers document that voting results respond to macro shocks, such as import competition (Autor et al., 2020; Dippel et al., 2022), economic crisis (Barros and Santos Silva, 2025), automation (Anelli et al., 2019), and austerity reforms (Fetzer, 2019). Similarly, at the individual level, many studies document that changes in economic conditions affect individuals’ policy preferences, including support for redistribution and welfare provision (e.g. Margalit, 2013; Martén, 2019; Ahlquist et al., 2020; Ballard-Rosa et al., 2021). In terms of political participation, the literature has focused mostly on turnout and has found mixed results, ranging from turnout increasing with local unemployment rates (Charles and Stephens, 2013; Burden and Wichowsky, 2014), decreasing after individual job loss (Emmenegger et al., 2017; Östermann and Lindgren, 2023), or not being affected at all by income changes (Jungkunz and Marx, 2022; Geys and Sørensen, 2025).
We contribute to this strand of literature by revealing that economic shocks not only affect the demand of voters, but also impact the supply side of politics. In particular, we document how layoffs increase individuals’ propensity to enter the political system, either as members of a political party or as candidates in local elections. In addition, our mass-layoff identification strategy, together with the size and granularity of our data, allows us to exploit exogenous changes in personal economic conditions and study the mechanisms behind individuals’ decisions to engage with politics after a layoff.
@@
This paper is structured as follows. Section 2 provides background information on party members and local council candidates in Brazil, and discusses theoretical predictions about the role of opportunity costs for political entry. Section 3 introduces the datasets used in the analyses. Section 4 presents the main results and heterogeneity analyses for the effect of job loss. In Section 5, we investigate the role of UI benefit eligibility. Finally, Section 6 examines potential implications for political selection, and Section 7 concludes.
@@ 2 Institutional and Theoretical Background
In Brazil, a significant portion of the population is engaged in politics. As mandated by the 1988 Brazilian constitution, voting is compulsory for all literate citizens aged 18 to 69.444The average turnout rate in Federal Elections in the last 20 years was about 80% (International Institute for Democracy and Electoral Assistance, 2024). Apart from voting, more than 16 million Brazilians were registered as members of a political party in 2020. This corresponds to about 11% of the electorate, among the highest shares observed across democratic countries. In addition, between 2000 and 2020, about 1.7 million individuals ran for office in local elections in Brazil. In this section, we describe the institutional environment in which individuals decide about their involvement as party members or local candidates and present a theoretical framework that illustrates the role of opportunity costs in this decision.
@@ 2.1 Party members
Membership recruitment in Brazil is predominantly a local phenomenon, fueled by internal competition at parties’ municipal conventions. Local elites, aiming to secure the party’s nomination, attract supporters by enlisting them in the party to build a loyal voting base at the convention and signal their potential as candidates (Sells, 2020; Frey, 2022). Supporters who become members of a winning political party are often rewarded with employment opportunities within the government bureaucracy (Brollo et al., 2017; Barbosa and Ferreira, 2023). As suggestive evidence of such benefits in our setting, Appendix Figure A.1 shows that laid-off individuals are more likely to start a public-sector job if they were member of a political party before the layoff, with no comparable increase observed for private-sector employment. The strategic importance of party members in local politics is further reflected in the cyclical pattern of new affiliations: as shown in Appendix Figure A.2, parties significantly expand their ranks in the year before local elections.